A jewelery maker pproduces necklaces that sell for $85 each. The jewelery maker's costs include $35 in materials for each necklace plus fixed costs of $1650. How many necklaces must the jeweler sell to break even?
What does the question mean?
Please show a step by step process.

Answer by richwmiller(17219)

1650=(85-35)*n
1650=(50)*n
1650/50=n
n=33